OpenStudy (anonymous):

There is an inverse relationship between x and y. If y is 20 when x is 2, what is x when y is 8?

OpenStudy (anonymous):

Let take some t. 20*t=8 then t=2/5 . t is the ratio or relationship in these to and since they y and x are inversely related use 1/t . Hence , 1/t = 5/2 and multiply it with 2 x =5

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

nvm thanks i got it